# Charge ordering and spontaneous topological Hall effect in bilayer skyrmion crystals

### v1

@@ Abstract
Motivated by these developments, we study in this Letter the impact of electron-electron interactions in zero field bilayer magnetic SkXs. Such bilayer magnetic SkXs with zero net topological charge and a vanishing THE may appear in centrosymmetric systems discussed above, or may be created in artificial heterostructures as depicted in Fig. 1(a). As a specific example, we study the effect of short-range interactions in a bilayer AF-SkX on the triangular lattice as depicted in Fig. 1(b) including both intralayer and interlayer interactions with Hartree-Fock theory. Our main finding is that interaction effects can induce various types of broken symmetry states. When the skyrmion bands are somewhat dispersive, moderate electron interactions lead to an electron crystal which may be viewed as an exciton condensate arising from neighboring Chern bands. This crystal does not break the SkX translation symmetry but breaks rotational and inversion symmetries of the SkX. For flatter Chern bands, even very weak interactions induce a weakly layer-polarized phase which also spontaneously breaks inversion symmetry. Both types of broken symmetry phases reveal a large THE. Our work suggests that it might be interesting to look for such spontaneous inversion symmetry breaking, perhaps via optical second harmonic generation accompanying the THE, in the zero field SkX phases of centrosymmetric systems including artificially engineered heterostructures.
Figure 1: Top: Schematic illustration of a heterostructure with inversion symmetry in the central layer which could lead to interfacial DM interactions of opposite sign at the two metal-magnet interfaces, which may be described by the Hamiltonian in Eq. 1, with the top and bottom layer skyrmions having opposite topological charge. Bottom: Specific bilayer Bloch skyrmion texture $\hat{{\cal S}}_{i,\ell}$ (as defined in Eq.(12)) on the triangular lattice chosen in our study has a 12-site unit cell, with the two layers having opposite ${\cal S}_{z}$ but the same ${\cal S}_{x},{\cal S}_{y}$ components.
@@ I Models of Bilayer skyrmions
where $c_{i,\ell,\sigma}$ represents an electron at site $i$ in layer $\ell$ with spin $\sigma$, and all the spin indices are summed over implicitly. The electrons in each layer have a nearest neighbor hopping term, and they are coupled via a local Hund’s electron-spin coupling to a SkX spin texture which is layer anti-ferromagnetic. The SkX spin configuration $\hat{{\cal S}}_{i,\ell}$ is plotted in Fig.1, where we have chosen ${{\cal S}}^{z}_{i,1}=-{{\cal S}}^{z}_{i,1}$, i.e. to be flipped between the layers, while ${{\cal S}}^{x}_{i,\ell}$ and ${{\cal S}}^{y}_{i,\ell}$ are chosen to be the same in each layer. Such a configuration ensures that the topological charge of the skyrmions in each layer is opposite. Further details of the SkX ansatz are given in Appendix A; see Eq. 12. This Hamiltonian describes an effective Kondo lattice between localized spins and conduction electrons. Such an antiferro-SkX states could potentially emerge in magnetic layers of a heterostructure stack, as depicted schematically in Fig. 1, and realized with micromagnetic simulations [33] and classical Monte Carlo results [60]. The Kondo coupling could then emerge through a proximity effect with a metallic layer in the heterostructure stack [53]. An alternative material realization would be in heavy fermion materials, including multipolar SkX configurations, induced by the RKKY interaction coupled to conduction electrons [52, 61]. Another material realization would be a layer-SkX state in TMD materials where the moiré potential provides the effective Hund’s coupling to the electrons [41]; this connection is further elaborated upon in the Discussion section.
@@ III Interaction induced electronic instabilities
We now turn to the role that the skyrmion crystal and electronic interactions have on charge ordering and band topology. To understand possible charge orders, we define the layer-resolved structure factor as
$n_{\ell}({{\bf{Q}}})={\int d^{2}\mathbf{r}}\>e^{i{{\bf{Q}}}\cdot\mathbf{r}}\expectationvalue{c^{\dagger}_{\mathbf{r},\ell,\sigma}c^{\phantom{\dagger}}_{\mathbf{r},\ell,\sigma}}\,.$ (6)
Using this, we can define the layer-combined charge and polarization at a given momentum as $n({{\bf{Q}}})=n_{1}({{\bf{Q}}})+n_{2}({{\bf{Q}}})$, and $\Delta n({{\bf{Q}}})=n_{1}({{\bf{Q}}})-n_{2}({{\bf{Q}}})$. Note that $n({{\bf{Q}}}=\Gamma)=2\bar{n}$ represents the combined uniform charge density per site, while $\Delta n({{\bf{Q}}}=\Gamma)$ represents the uniform polarization per site. For further information about each phase, The connected part of the structure factors can also demonstrate peaks at non-zero ${{\bf{Q}}}$. These peaks would indicate additional charge ordering or modulated polarization.
The bare SkX state provides a small amount of charge redistribution, which is documented in Fig. 5. It is clear from the peak at ${{\bf{Q}}}=0$ that most of the charge is distributed uniformly. However, from the connected structure factor, we see there are small but non-zero charge distribution at other momenta like the $M$ points, as well as the SkX ordering vector, which are the $K/2$ points.
Figure 5: The charge distribution in momentum space in the presence of the skyrmion at $J_{h}/t=4$ when the $13^{\text{th}}$ degenerate band is half-filled at $\bar{n}=25/48$. (a) The absolute value of the total $|n({{\bf{Q}}})|$ in the triangular Brillouin zone. (b) Only the connected piece of part (a). The green hexagon marks the triangular Brillouin zone, while the smaller yellow one is that of the SkX unit cell. The $\times$ points mark the reciprocal lattice sites corresponding to the SkX unit cell.
@@
Figure 6: For $J_{h}/t=4$, The topological response $\sigma_{xy}$, the polarization $\delta n$, and the static structure factor $\Delta n(Q)$ of the polarization that emerge as a continuous transition with increasing the interlayer repulsion, $U$. Here we consider $V=0$. The underlying skyrmion ordering vector is ${{\bf{Q}}}={{\bf{K}}}/2$.
@@
Fig. 6 demonstrates a net layer polarization due to splitting of the degenerate bands. Subtracting this disconnected piece demonstrates residual charge distribution at incommensurate ${{\bf{Q}}}$, proportional to the SkX as shown in Fig. 13 in Appendix. C. As such, one band becomes fully filled and the system exhibits a quantized Hall response, a spontaneous topological Hall effect. The polarization order occurs at ${{\bf{Q}}}=\Gamma$, with uniform polarization in one layer.
@@
Figure 7: The structure factor $\Delta n$ and topological Hall response $\sigma_{xy}$ as a function of $U$ and $V$ for the flat-band under discussion for $J_{h}=4$.
@@
At $J_{h}/t=2$, the $13^{\text{th}}$ band is not isolated and the system is metallic. For half-filling this band ($\bar{n}=25/48$), the system is stable to layer polarization for small $U$, $(U\lesssim 2)$. However, for moderate $V/t\sim 1.5$, there is a rotational symmetry breaking transition ($C_{6z}\rightarrow C_{3z}$), demonstrated in Fig. 8 and Fig. 9. Here instead it is the breaking of $C_{2z}$ which breaks the centrosymmetry ${\mathcal{C}}$ and leads to a spontaneous topological Hall effect. The structure factor in Fig. 14 in Appendix C demonstrates this rotational symmetry breaking. Throughout the ordered phase, there is a considerable unquantized Hall response $\sigma_{xy}$. The large number of bands intersecting prevents any quantitative conclusions to be drawn from precise values of $\sigma_{xy}$. This transition occurs due to band-touching and reformation points rather than a simple degeneracy lifting that occurs for larger $J_{h}$. This phenomenology resembles Landau level mixing. At large $J_{h}$, the band gaps are too large and the Fermi surface is too featureless for any weak-coupling instabilities to crystalline symmetry broken states.
Figure 9: For $J_{h}/t=2$, The topological response $\sigma_{xy}$, the polarization $\Delta n$, and the static structure factor $\Delta n({{\bf{Q}}})$ of the charge-order that emerge as a continuous transition with increasing the intralayer repulsion, $V$. the ordering vector is ${{\bf{Q}}}={{\bf{K}}}$. The data is plotted for $U/t=0.6$ as a representative sample.
@@
Figure 10: The structure factor $\Delta n({{\bf{Q}}})$ and topological Hall response $\sigma_{xy}$ as a function of $U$ and $V$ for the band under discussion for $J_{h}/t=2$. Here, there is no weak-coupling limit at ${{\bf{Q}}}=\bm{\Gamma}$. Rather there is a first-order transition to charge ordering at the SkX unit vector ${{\bf{Q}}}={{\bf{K}}}/2$, with a real space representation in Fig. 8.
@@ IV Discussion
This work demonstrates several intriguing properties of skyrmion crystals and points to future directions. We demonstrate that there are flat bands with a propensity to form crystalline analogs of Landau levels within the spectrum of SkX. We focus on a particular flat-band, just above half-filling, and explore the symmetry broken instabilities at partial filling of this degenerate band. A considerable amount of research has focused on fine-tuned or “magic-angle” moiré configurations to realize flat bands [78]. There is however an increased realization that magnetic order can result in similar flat bands [53]. This offers intriguing experimental directions to explore in magnetic bilayer systems.
@@
Future theoretical directions can focus on states within these bands with genuine topological order, which will require numerical analysis beyond the mean-field level. Of particular interest in future studies is the required proximity to Landau level analogs for SkX flat bands to exhibit fractionalization [64, 59, 54]. SkX with $Q>1$ may also in turn host bands with $C>1$. If these bands allow for FCI states, there may be many qualitative and quantitative differences to previous realizations [79].
Within Hartree-Fock however, these states already offer interesting broken-symmetry instabilities. We find at small $J_{h}$, intra-layer repulsion $V$ can form crystalline formation with nonzero Hall response. One potential direction would be to explore what interactions would be necessary to realize “anomalous Hall crystals” where the resulting crystalline state has an isolated band with non-zero Chern number $C$ [80, 81]. At larger $J_{h}$, combinations of interlayer repulsion $U$ and $V$ compete to drive an inversion-breaking polarization transition. This phenomenology and competition can shed light on when polarized states are necessary to realize non-zero Hall effects [67, 68, 69, 70, 71, 72]. There are intriguing possibilities of coherent (rather than polarized) states that still exhibit an AHE [82]. Theoretical and experimental work on realizing such states in AF-SkX states would be enlightening. The ability to control these effects, through interlayer spacing for example, may offer control for topological and spintronics applications seeking to harness SkX states for quantum information storage and manipulation [6, 4, 7].
@@ Acknowledgements.
We thank D. Guerci, J. Huxford, and N. Paul for insightful discussions. We acknowledge support from the Natural Sciences and Engineering Research Council (NSERC) of Canada. A.H. acknowledges support from a NSERC Graduate Fellowship (PGS-D). A.H. acknowledges support from a pre-doctoral fellowship at the Flatiron Institute while part of this work was being completed. The Flatiron Institute is a division of the Simons Foundation. T.G. acknowledges support from the Center For Quantum Information and Quantum Control (CQIQC) Summer Fellowship. Numerical computations were performed on the Niagara supercomputer at the SciNet HPC Consortium and the Digital Research Alliance of Canada.
@@ Appendix A Skyrmion Classification
Classification of skyrmion Crystal phases are determined by their homotopy group $\pi_{2}(\text{SO}(3))$. This topological charge is constructed from the combination of the skyrmion winding number and polarity $Q=P\cdot\mathcal{W}$. For a spin-texture given by $\hat{\mathcal{S}}^{\ell}_{\alpha}(\mathbf{r})$ where $\alpha\in\{x,y,z\}$ and $\ell\in\{1,2\}$ represents the layer, we have
@@
[…]
@@
where $x\,,y\,,$ and $r$ are measured with respect to the SkX unit-cell origin. Rotating the spins by $\pi/2$ in plane switches from Bloch to Neel skyrmions. Lastly, we are also agnostic to the helicity of the SkX phases (Neel or Bloch as shown in Fig. 1) because a spin-rotation can convert between the two and as such they do not affect band-structure results.

### v2

@@ Abstract
Motivated by these developments, we study in this Letter the impact of electron-electron interactions in zero field bilayer magnetic SkXs. Such bilayer magnetic SkXs with zero net topological charge and a vanishing THE may appear in centrosymmetric systems discussed above, or may be created in artificial heterostructures as depicted in Fig. 1(a). As a specific example, we consider a bilayer SkX on the triangular lattice as depicted in Fig. 1(b), where $S_{z}$ is flipped between layer-1 and layer-2 while $S_{x},S_{y}$ remain parallel so that the topological charge $\sim\int\mathbf{S}\cdot\partial_{x}\mathbf{S}\times\partial_{y}\mathbf{S}$ is opposite on the two layers. Our focus is on the impact of electron-electron interactions on such SkXs, including the impact of both intralayer and interlayer interactions with Hartree-Fock theory. Our main finding is that interaction effects can induce various types of broken symmetry states in such SkX states. When the skyrmion bands are somewhat dispersive, moderate electron interactions lead to an electron crystal which may be viewed as an exciton condensate arising from neighboring Chern bands. This crystal does not break the SkX translation symmetry but breaks rotational and inversion symmetries of the SkX. For flatter Chern bands, even very weak interactions induce a weakly layer-polarized phase which also spontaneously breaks inversion symmetry. Both types of broken symmetry phases reveal a large THE. Our work suggests that it might be interesting to look for such spontaneous inversion symmetry breaking, perhaps via optical second harmonic generation accompanying the THE, in the zero field SkX phases of centrosymmetric systems including artificially engineered heterostructures.
Figure 1: Top: Schematic illustration of a heterostructure with inversion symmetry in the central layer which could lead to interfacial DM interactions of opposite sign at the two metal-magnet interfaces, which may be described by the Hamiltonian in Eq. 1, with the top and bottom layer Néel skyrmions having opposite topological charge. Our key results also remain valid for Bloch skyrmions. Bottom: Example of a bilayer Bloch skyrmion texture $\hat{{\cal S}}_{i,\ell}$ (as defined in Eq.(12)) on the triangular lattice (with 12-site unit cell in our work), and the two layers having opposite ${\cal S}_{z}$ but the same ${\cal S}_{x},{\cal S}_{y}$.
@@ I Models of Bilayer skyrmions
where $c_{i,\ell,\sigma}$ represents an electron at site $i$ in layer $\ell$ with spin $\sigma$, and all the spin indices are summed over implicitly. The electrons in each layer have a nearest neighbor hopping term, and they are coupled via a local Hund’s electron-spin coupling to a SkX spin texture which is layer anti-ferromagnetic. The SkX spin configuration $\hat{{\cal S}}_{i,\ell}$ is plotted in Fig.1, where we have chosen ${{\cal S}}^{z}_{i,2}=-{{\cal S}}^{z}_{i,1}$, i.e. to be flipped between the layers, while ${{\cal S}}^{x}_{i,\ell}$ and ${{\cal S}}^{y}_{i,\ell}$ are chosen to be the same in each layer. Such a configuration ensures that the topological charge of the skyrmions in each layer is opposite. Further details of the SkX ansatz are given in Appendix A; see Eq. 12. This Hamiltonian describes an effective Kondo lattice between localized spins and conduction electrons. Such an antiferro-SkX states could potentially emerge in magnetic layers of a heterostructure stack, as depicted schematically in Fig. 1, and realized with micromagnetic simulations [33] and classical Monte Carlo results [60]. The Kondo coupling could then emerge through a proximity effect with a metallic layer in the heterostructure stack [53]. An alternative material realization would be in heavy fermion materials, including multipolar SkX configurations, induced by the RKKY interaction coupled to conduction electrons [52, 61]. Another material realization would be a layer-SkX state in TMD materials where the moiré potential provides the effective Hund’s coupling to the electrons [41]; this connection is further elaborated upon in the Discussion section.
@@ III Interaction induced electronic instabilities
We now turn to the role that the skyrmion crystal and electronic interactions have on charge ordering and band topology. To understand possible charge orders, we define the layer-resolved charge density in momentum space as
$n_{\ell}({{\bf{Q}}})=\sum_{\sigma}{\int d^{2}\mathbf{r}}\>e^{i{{\bf{Q}}}\cdot\mathbf{r}}\expectationvalue{c^{\dagger}_{\mathbf{r},\ell,\sigma}c^{\phantom{\dagger}}_{\mathbf{r},\ell,\sigma}}\,.$ (6)
Using this, we can define the layer-combined charge and polarization at a given momentum as $n({{\bf{Q}}})=n_{1}({{\bf{Q}}})+n_{2}({{\bf{Q}}})$, and $\Delta n({{\bf{Q}}})=n_{1}({{\bf{Q}}})-n_{2}({{\bf{Q}}})$. Note that $n({{\bf{Q}}}=\Gamma)=2\bar{n}$ represents the combined uniform charge density per site, while $\Delta n({{\bf{Q}}}=\Gamma)$ represents the uniform polarization per site, where $\Gamma=(0,0)$. To study the non-uniform part of the charge distribution, we also plot the Fourier transform of the charge modulation, $n_{\ell}({{\bf{Q}}})-\overline{n}_{\ell}$. These peaks would indicate additional charge ordering or modulated polarization.
The bare SkX state provides a small amount of charge redistribution, which is documented in Fig. 5. It is clear from the peak at ${{\bf{Q}}}=0$ that most of the charge is distributed uniformly. However, from the modulated charge distribution, we see there are small but non-zero charge distribution at other momenta like the $M$ points, as well as the SkX ordering vector, which are the $K/2$ points.
Figure 5: The charge distribution in momentum space in the presence of the skyrmion at $J_{h}/t=4$ when the $13^{\text{th}}$ degenerate band is half-filled at $\bar{n}=25/48$. (a) The absolute value of the total $|n({{\bf{Q}}})|$ in the triangular Brillouin zone. (b) The finite $q$ piece of part (a). The green hexagon marks the triangular Brillouin zone, while the smaller yellow one is that of the SkX unit cell. The $\times$ points mark the reciprocal lattice sites corresponding to the SkX unit cell.
@@
Figure 6: For $J_{h}/t=4$, The topological Hall response $\sigma_{xy}$, the polarization $\Delta n$, and the charge order $\Delta n(Q)$ that emerge as a continuous transition with increasing the interlayer repulsion, $U$. Here we consider $V=0$. The underlying skyrmion ordering vector is ${{\bf{Q}}}={{\bf{K}}}/2$.
@@
Fig. 6 demonstrates a net layer polarization due to splitting of the degenerate bands. Subtracting this uniform piece demonstrates residual charge distribution at incommensurate $\bf{Q}$}}, proportional to the SkX as shown in Fig. 13 in Appendix. C. As such, one band becomes fully filled and the system exhibits a quantized Hall response, a spontaneous topological Hall effect. The polarization order occurs at ${{\bf{Q}}}=\Gamma$, with uniform polarization in one layer.
@@
Figure 7: The polarization $\Delta n$ and topological Hall response $\sigma_{xy}$ as a function of $U$ and $V$ for the flat-band under discussion for $J_{h}=4$.
@@
At $J_{h}/t=2$, the $13^{\text{th}}$ band is not isolated and the system is metallic. For half-filling this band ($\bar{n}=25/48$), the system is stable to layer polarization for small $U$, $(U\lesssim 2)$. However, for moderate $V/t\sim 1.5$, there is a rotational symmetry breaking transition ($C_{6z}\rightarrow C_{3z}$), demonstrated in Fig. 8 and Fig. 9. Here instead it is the breaking of $C_{2z}$ which breaks the centrosymmetry $\mathcal{C}$} and leads to a spontaneous topological Hall effect. The charge modulation in Fig. C in Appendix C demonstrates this rotational symmetry breaking. Throughout the ordered phase, there is a considerable unquantized Hall response $\sigma_{xy}$. The large number of bands intersecting prevents any quantitative conclusions to be drawn from precise values of $\sigma_{xy}$. This transition occurs due to band-touching and reformation points rather than a simple degeneracy lifting that occurs for larger $J_{h}$. This phenomenology resembles Landau level mixing. At large $J_{h}$, the band gaps are too large and the Fermi surface is too featureless for any weak-coupling instabilities to crystalline symmetry broken states.
Figure 9: For $J_{h}/t=2$, The topological Hall response $\sigma_{xy}$, the polarization $\Delta n$, and the charge order $\Delta n({{\bf{Q}}})$ that emerge as a continuous transition with increasing the intralayer repulsion, $V$. the ordering vector is ${{\bf{Q}}}={{\bf{K}}}$. The data is plotted for $U/t=0.6$ as a representative sample.
@@
Figure 10: The charge order $\Delta n({{\bf{Q}}})$ and topological Hall response $\sigma_{xy}$ as a function of $U$ and $V$ for the band under discussion for $J_{h}/t=2$. Here, there is no weak-coupling limit at ${{\bf{Q}}}=\bm{\Gamma}$. Rather there is a first-order transition to charge ordering at the SkX unit vector ${{\bf{Q}}}={{\bf{K}}}/2$, with a real space representation in Fig. 8.
@@ IV Discussion
This work demonstrates several intriguing properties of skyrmion crystals and points to future directions. We demonstrate that there are flat bands with a propensity to form crystalline analogs of Landau levels within the spectrum of SkX. We focus on a particular flat-band, just above half-filling, and explore the symmetry broken instabilities at partial filling of this degenerate band.
For large $J_{h}/t$, the polarization is suppressed by the large band-gaps, similar to the suppression of Landau level mixing for small interaction strength. For small $J_{h}/t$, there is the possibility of the nearest neighbor $V$ term driving band-mixings and reordering the charges. In contrast, the large density of states of the flatter bands at large $J_{h}/t$ allows for weak-coupling instabilities to layer polarization without any single layer band-mixing.
A considerable amount of research has focused on fine-tuned or “magic-angle” moiré configurations to realize flat bands [78]. There is however an increased realization that magnetic order can result in similar flat bands [53]. This offers intriguing experimental directions to explore in magnetic bilayer systems.
@@
Future theoretical directions can focus on states within these bands with genuine topological order, which will require numerical analysis beyond the mean-field level. Of particular interest in future studies is the required proximity to Landau level analogs for SkX flat bands to exhibit fractionalization [64, 59, 54]. SkX consisting of SU(N>1)(N>1) spinors or with topological charge $Q>1$ may also in turn host bands with $C>1$. If these bands allow for FCI states, there may be many qualitative and quantitative differences to previous realizations [79].
Within Hartree-Fock however, these states already offer interesting broken-symmetry instabilities. We find at small $J_{h}$, intra-layer repulsion $V$ can form crystalline formation with nonzero Hall response. One potential direction would be to explore what interactions would be necessary to realize “anomalous Hall crystals” where the resulting crystalline state has an isolated band with non-zero Chern number $C$ [80, 81, 82]. At larger $J_{h}$, combinations of interlayer repulsion $U$ and $V$ compete to drive an inversion-breaking polarization transition. This phenomenology and competition can shed light on when polarized states are necessary to realize non-zero Hall effects [67, 68, 69, 70, 71, 72]. There are intriguing possibilities of coherent (rather than polarized) states that still exhibit an AHE [83]. Theoretical and experimental work on realizing such states in AF-SkX states would be enlightening. The ability to control these effects, through interlayer spacing for example, may offer control for topological and spintronics applications seeking to harness SkX states for quantum information storage and manipulation [6, 4, 7].
@@ Acknowledgements.
We thank D. Guerci, J. Huxford, N. Paul, and A. Reddy for insightful discussions. We acknowledge support from the Natural Sciences and Engineering Research Council (NSERC) of Canada. A.H. acknowledges support from a NSERC Graduate Fellowship (PGS-D). A.H. acknowledges support from a pre-doctoral fellowship at the Flatiron Institute while part of this work was being completed. The Flatiron Institute is a division of the Simons Foundation. T.G. acknowledges support from the Center For Quantum Information and Quantum Control (CQIQC) Summer Fellowship. Numerical computations were performed on the Niagara supercomputer at the SciNet HPC Consortium and the Digital Research Alliance of Canada.
@@ Appendix A Skyrmion Classification
Classification of skyrmion Crystal phases are determined by their homotopy group $\pi_{3}(\text{SO}(3))$. This topological charge is constructed from the combination of the skyrmion winding number and polarity $Q=P\cdot\mathcal{W}$. For a spin-texture given by $\hat{\mathcal{S}}^{\ell}_{\alpha}(\mathbf{r})$ where $\alpha\in\{x,y,z\}$ and $\ell\in\{1,2\}$ represents the layer, we have
@@
$\begin{split}Q^{\ell}&=\frac{1}{4\pi}\int_{0}^{\infty}dr\sin\Theta^{\ell}\frac{d\Theta^{\ell}}{dr}\cdot\int_{0}^{2\pi}d\phi\frac{d\Phi^{\ell}}{d\phi}\\ &=\frac{\left[\hat{{\cal S}}_{z}^{\ell}(0)-\hat{{\cal S}}_{z}^{\ell}(\infty)\right]}{2}\frac{[\Phi^{\ell}(2\pi)-\Phi^{\ell}(0)]}{2\pi}=P^{\ell}\cdot\mathcal{W}^{\ell}\end{split}$ (11)
@@
where x,y,x\,,y\,, and $r$ are measured with respect to the SkX unit-cell origin. Rotating the spins by $\pi/2$ in plane switches from Bloch to Neel skyrmions. Lastly, we are also agnostic to the helicity of the SkX phases (Neel or Bloch as shown in Fig. 1) because a spin-rotation can convert between the two and as such they do not affect band-structure results.